Next one-year period, forecasted inflation rate is 5% in China and 3% in US. Current spot exchange rate is RMB4.00/USD (RMB is the name of Chinese currency). What is the forecasted spot exchange rate in one year according to relative PPP?

A.

RMB4.08/USD

B.

RMB4.12/USD

C.

RMB4.27/USD

D.

RMB4.40/USD
Answer #1

Hi,

Here current spot rate = RMB 4/USD

China inflation rate = 5%

US inflation rate =3%

US -China inflation rate = 5-3 =2%

Hence as per relative PPP

forecasted spot exchange rate = (1+China inflation rate - US inflation rate)*current spot rate

=(1+2%)*4

=4*1.02 = RMB 4.08/USD

Hence option A is correct here

Thanks